# Self-propulsion of floating ice blocks caused by melting in water

### v2

@@ Abstract
We show that floating ice blocks with asymmetric shapes can self-propel with significant speeds due to buoyancy driven currents caused by melting. Model right-angle ice wedges are found to move in the direction opposite to the gravity current, which descends along the longest inclined side, in water baths with temperatures between $10^{\circ}$ and $30^{\circ}$. We describe the measured speed as a function of the length and angle of the inclined side, and the temperature of the bath in terms of a propulsion model which incorporates the cooling of the surrounding fluid by the melting of ice. The heat pulled from the surrounding liquid by the melting ice block is shown to lead to propulsion which is balanced by drag. We further show that the ice block moves robustly in a saltwater bath with ocean-like salinity and maintains the same direction of motion as in freshwater. A simplified model is further developed to describe the propulsion of asymmetric ice blocks in saltwater, incorporating the effects of rising meltwater and the sinking of the surrounding bath water due to cooling. For sufficiently large temperature, we find that the cooling-induced sinking flow generates a stronger force than the upward flow from the meltwater. Consequently, the net propulsion force is in the same direction and nearly the same magnitude as that observed in freshwater. These findings suggest that melting-driven propulsion may be relevant to the motion of icebergs in sufficiently warm oceanic environments.
@@ I Introduction
An additional source of energy is not required to create a heat flux and convection current for a melting block floating in a bath at a temperature different from the melting temperature. Previously, Dorbolo, et al. [8] related the spinning of floating ice disks to the convection flow driven by melting. No translation was reported because the disks were symmetric and were constrained to rotate by fixing the center position using magnets. Recently, it was demonstrated that a boat incorporating an inclined solute material like salt or sugar can propel rapidly due to the solutal convection flow driven by the dissolution [9]. We build on that study by investigating the case of asymmetric ice blocks as they melt in warm water. As in the case of dissolution, the convection flow generated by density variations due to a phase change from solid to liquid leads to self-propulsion, if the convection flow displays a forward-backward asymmetry. The flow then has a non-zero horizontal momentum component, which results in an oppositely directed reaction propulsion force on the body, that can be estimated using a momentum balance. However, melting-driven propulsion is not equivalent to dissolution-driven propulsion are fundamentally different even though both are phase changes from solid to liquid [10].
[…]
@@
Here, we investigate the kinematics of ice blocks which have asymmetric shapes while floating in a water bath and show that they can not only rotate but translate with significant speeds. We find a typical propulsion velocity of about 3 mm s-1 for triangular ice prisms with an inclined long side of approximately 20 cm and width of approximately 10 cm, floating in a water bath held at a temperature of about $T_{b}=22\,^{\circ}$, comparable to those observed with dissolving bodies. After presenting the experimental methods in Section II, we demonstrate and quantify the melting-driven propulsion mechanism in Section III. We use shadowgraph imaging [19, 20] to simultaneously track the motion of the block and visualize the buoyancy convection flow. A phenomenological model relating the melting velocity to the terminal speed is developed in Section IV to explain the magnitude of the observed ice block translation velocities as a function of their size, inclination, and bath temperature. In the limit where the heat required to raise the temperature of the ice block to the melting temperature is relatively small, our model finds that although the latent heat plays an important role in the dynamics and determines the time over which the block melts, its actual magnitude does not significantly affect the propulsion speed. In Section IV.4, we compare the measured values over a large range of parameters with the predictions of the model. Then, we demonstrate with experiments that the melting-driven propulsion mechanism extends to baths with ocean water salinity in Section V, and adapt our model to take into account the rising fresh water flow and the falling convection flow due to the cooling of the salt bath. We address the possible relevance of the melting-driven propulsion mechanism to icebergs in oceans in Section VI, before summarizing our main results in Section VII. Although our results cannot be directly applied to icebergs under typical conditions, we argue that this effect may contribute to the motion of icebergs drifting in warmer subpolar regions.
@@
Figure 1: (a) A schematic of a melting ice block floating in a water bath. The block is a right-angled triangle prism. The hypotenuse has a length $L$ and is inclined at an angle $\theta$ relatively to the horizontal. The ice block propels along the horizontal coordinate $x$ with a velocity $\bm{U_{b}}$ in steady state. (b) A side view image of a right-angle ice block ($L_{h}=163$ mm, $W=124$ mm, $H=65$ mm, $\theta=19.5^{\circ}$, $T_{b}=22\,^{\circ}$). (c) A profile view image of the floating ice block with a parabolic mirror placed behind the tank containing freshwater. The dense cold water leads to convection flow which is visible below the ice block. The mirror is then used for shadowgraph imaging. (d) A shadowgraph picture shows the ice block moving to the right with the convection directed toward the rear of the block (also see Movie S1 [21]).
@@ II Experimental methods
We cast asymmetric ice blocks with a simple model geometry: a right-angled triangle prism. The ice block has a hypotenuse of length $L$ and is inclined at an angle $\theta$ with respect to the horizontal and is shown schematically in Fig. 1(a). The ice blocks are cast in molds with horizontal length $L_{h}$, width $W$, and side $H$ filled with filtered demineralized water and freezing at $-15\,^{\circ}$. Approximately one hundred blocks were cast with various sizes ranging over a few tens of centimeters. The ice blocks contain bubbles which typically results in about 10% of air by volume, similar to icebergs [22]. A few clear rectangular ice blocks which are nearly transparent and defect-free were obtained from the Nice Company (https://www.thenicecompanyparis.com/fr), and cut into triangular prisms with a hot wire. Experiments performed with the clear ice blocks were found to give the same results as the ice blocks with trapped bubbles. The data sets corresponding to various blocks shapes and measurement protocols are listed in Appendix A.
@@
Prior to the commencement of an experiment, an ice block is left to rest at room temperature of about $20\,^{\circ}$ for about ten minutes in order to avoid thermal shock when immersing it into the bath. During this time, the ice temperature approaches the melting temperature of ice. Then, the block is carefully placed with its right angle on the top in a water bath with dimensions that are large compared to the block size. Flotation equilibrium corresponds to the vertical alignment of the gravity center and the center of the immersed part and is typically reached after few oscillations over a few seconds. A side view image of a floating ice block is shown in Fig. 1(b). The actual hypotenuse inclination angle differs from $\theta$ by few degrees due to the buoyancy equilibrium of the asymmetric block (see Appendix B). Yet, we ignore this difference since it appears to have little effect on the measured trends. Ice blocks with $\theta>39.6^{\circ}$ are gravitationally unstable, and we thus focus our study to ice blocks with $\theta$ below this angle.
@@
The ice blocks and their motion in the bath are either observed with a camera from the side using a shadowgraph imaging [19, 20], or from the top. To observe the dynamics with shadowgraphy, we use a glass tank with dimensions $116\times 46$ cm2 filled up to $24$ cm with tap water corresponding to a volume of about $128$ liters. For most experiments the bath temperature $T_{b}$ is the ambient temperature. However, for a few experiments $T_{b}$ has been varied between $10.4$ and $30\,^{\circ}$ using a cooling / heating circulator. A parabolic mirror with diameter $406$ mm and focal length $1800$ mm is placed on one side of the glass tank at a distance of about 100 mm and oriented parallel to it, as depicted in Fig. 1(c). A small Light Emitting Diode (LED) located at the focus of the parabolic mirror is used for illumination, and a digital camera is located at focus of the mirror by the means of a semi-reflective plate. The resulting light beam with nearly parallel rays is refracted by the variations of optical indices due to temperature variations. The projected image captured by the digital camera integrates the density variations along the width of the tank. For a bath of fresh water whose temperature is above maximum density of water at $T_{c}=3.98\,^{\circ}$, the colder convection plumes sink and and refract ambient light. The plumes appear more clearly in the image shown in Fig. 1(d) generated with shadowgraphy. A small black circle marks the center of the mirror and is visible on some shadowgraph images and in the movies [21]. In these experiments, two 1 mm in diameter nylon wires are positioned just below the surface to guide the motion of the ice block, separated by a distance slightly greater than $W$. This helps maintain the distance between the block and the camera and limits rotation, facilitating observations. In the experiments where the ice blocks are viewed from the top, we use a glass tank with dimensions $90.5\times 44.5$ cm2, filled up to a height of at least $25$ cm with filtered water corresponding to a volume of about $100$ liters. The ice block is free to move in any horizontal direction and data is taken while the block is away from the tank sidewalls. This can be used to study the relative stability of the ice block trajectory.
@@ III Experimental demonstration of melting-driven propulsion
Figure 2: (a) A view of an ice block floating at the water surface while moving in the direction of the observer after being immersed in the bath for about 8 minutes. A groove pattern generated by melting appears on the bottom inclined face. (b) A schematic cross section of the block illustrating the length-wise grooves. This melting pattern is carved by the backward descending flow. (c) An ice block with bottom surface facing viewer after it is removed from the water bath following approximately 11 minutes of immersion. Typical groove width and depth are about $20$ mm and $5$ mm, respectively. This example corresponds to Data Set L. ($L=176$ mm, $\theta=19.5^{\circ}$, $T_{b}=22.1\,^{\circ}$).
@@
Figure 3: (a) Right-angle triangle horizontal ice block position $X_{B}$ as a function of time $t$ plotted along with a linear fit (dashed line) corresponding to a slope $U_{b}=3.02$ mm s-1. The velocity of the block $U(t)$ obtained from $X_{B}$. This experiment corresponds to the block depicted in Movie S2 [21], where the block is kicked manually in the direction opposite to the motion induced by propulsion at $t=8$ s. After a transient, the block recovers the same velocity (dashed line).
Figure 4: (a) For a symmetric rectangular ice block position $X_{B}$ shows the slow drift of the block over time (also see Movie S3 [21]). The block dimensions: $100\times 40\times 40$ mm3, $\theta=0^{\circ}$, $T_{b}=20.1\,^{\circ}$. (d) The corresponding velocity (averaged over $10$ s) shows sudden changes in velocity due to capsizing events at $t=210$ s and $t=506$ s.
@@
To illustrate that the forward motion is quite robust, we manually kicked a block moving at constant velocity in the direction opposite to its motion at approximately 5 s (see Movie S2 [21]). The measured ice block position $X_{B}$, obtained by processing the shadowgraph images, is plotted as a function of time $t$ in Fig. 3(a). The instantaneous velocity $U(t)$ is computed over a moving $1$ s time interval and plotted in Fig. 3(b). The terminal velocity $U_{b}$ is obtained by fitting a line to Fig. 3(a) once steady motion is resumed. We plot a dashed horizontal time in Fig. 3(b) and observe that the block accelerates and reaches the pre-kick velocity $U_{b}\approx 3.02$ mm s-1 in about 30 s. By contrast, symmetric rectangular ice blocks do not show net translation motion (see Movie S3 [21]). We plot $X_{B}(t)$ and $U(t)$ for this symmetric ice block in Fig. 4(a), and Fig. 4(b), respectively. The time scales for motion are significantly larger compared with those observed in the asymmetric examples. The ice block motion shows no clear direction, and capsizing events occur suddenly and modify the convection flow that can change the direction of the ice block drift.
@@
We also perform experiments with ice blocks which are unconstrained laterally and we observe their motion in the tank from above. While the sides of the block melt on average at a rate of a few tens of microns per second, we observe that the block inclination is approximately conserved, at least during the first ten minutes or so, even as the ice block shrinks and the edges round over time. However, we find that the movement is not exactly rectilinear, but can show rotation in addition to translation. This rotation may be due to lateral asymmetries in the ice block, and/or caused by the destabilization of the falling convection flow into a vortex as reported with spinning ice disks [8, 27]. The rotational motion becomes increasingly important over time, as the ice block decreases in size and its shape deviates significantly from being prismatic as it completely melts. For this reason, systematic data is taken during the first ten minutes while the overall ice block shape appears similar as the one initially cast, barring the grooves. In these conditions, the dimensions of the blocks considered are determined at the start of the run. The different protocols with various ice blocks shapes are summarized in Appendix A.
@@
The measured $U_{b}$ show more significant experimental variability compared with dissolution propelled bodies [9] that can be attributed to several factors. The density contrast driving the convection flow is stronger by two orders of magnitude in the case of sugar or salt dissolution compared with the density contrast observed in ice blocks when $T_{b}=20\,^{\circ}$. The typical scale of the plumes about $3\delta_{i}\approx 5$ mm as discussed in Section IV.1 is larger compared to the block size leading to greater fluctuations in their trajectories. The shapes of the ice blocks are also less ideal and show around 5% variations. The ice blocks also intermittently release trapped bubbles during melting that further perturb transiently the flow below the ice block. Recent experiments [28] report an enhanced melting rate for vertical ice walls, due to the additional buoyancy associated with the released bubbles. However, in our experiments, we do not find significant differences on average for $U_{b}$ between the standard ice blocks and those made of clear ice without bubbles. Notwithstanding the fluctuations, we observe a robust directed motion with velocity $U_{b}\approx 3$ mm s-1 for an ice block with $L=20$ cm melting in a bath at $\,{}^{\circ}$ which can be easily noted with the naked eye. The corresponding Reynolds number $Re=\frac{L\,U_{b}}{\nu}$ is of order $600$, with the kinematic viscosity of water $\nu\approx 10^{-6}$ m2 s-1. Thus, the fluid flow is in the inertial regime, and we propose a phenomenological model based on momentum balance inspired by Chaigne, et al. [9].

### v3

@@ Abstract
We show that floating ice blocks with asymmetric shapes can self-propel with significant speeds due to buoyancy driven currents caused by melting. In water baths with temperatures between $10\,^{\circ}$C and $30\,^{\circ}$C, model right-angle ice wedges are found to move in the direction opposite to the gravity current which descends along the longest inclined side. We describe the measured speed as a function of the length and angle of the inclined side, and the temperature of the bath in terms of a propulsion model which incorporates the cooling of the surrounding fluid by the melting of ice. The heat pulled from the surrounding liquid by the melting ice block generates a thermal convection flow, leading to momentum exchange and a net propulsion force. The translation velocity is explained by balancing the propulsion force by drag. We further show that the ice block moves robustly in a saltwater bath with ocean-like salinity and maintains the same direction of motion as in freshwater. A simplified model is further developed to describe the propulsion of asymmetric ice blocks in saltwater, incorporating the effects of rising meltwater and the sinking of the surrounding bath water due to cooling. For sufficiently large temperature, we find that the cooling-induced sinking flow generates a stronger force than the upward flow from the meltwater. Consequently, the net propulsion force is in the same direction and nearly the same magnitude as that observed in freshwater. These findings suggest that melting-driven propulsion may be relevant to the motion of icebergs in sufficiently warm oceanic environments.
@@ I Introduction
An additional source of energy is not required to create a heat flux and convection current for a melting block floating in a bath at a temperature different from the melting temperature. Previously, Dorbolo, et al. [8] related the spinning of floating ice disks to the convection flow driven by melting. No translation was reported because the disks were symmetric and were constrained to rotate by fixing the center position using magnets. Recently, it was demonstrated that a boat incorporating an inclined solute material like salt or sugar can propel rapidly due to the solutal convection flow driven by the dissolution [9]. We build on that study by investigating the case of asymmetric ice blocks as they melt in warm water. As in the case of dissolution, the convection flow generated by density variations due to a phase change from solid to liquid leads to self-propulsion, if the convection flow displays a forward-backward asymmetry. The flow then has a non-zero horizontal momentum component, which results in an oppositely directed reaction propulsion force on the body, that can be estimated using a momentum balance. However, melting-driven propulsion is not equivalent to dissolution-driven propulsion. Melting and dissolution are fundamentally different even though both are phase changes from solid to liquid [10].
Melting is driven by temperature transport and the Stefan condition at the solid-liquid interface results from an energy balance involving latent heat. By contrast, dissolution of salt or sugar in water is driven by solute transport and the boundary condition corresponds to the solute mass conservation balance. Moreover, in the case of ice melting, the density of liquid water has a non-monotonic evolution close to the melting temperature with a maximal density at the temperature of maximum density $T_{c}$ equal to $3.98\,^{\circ}$C in fresh water, whereas the density increases linearly with solute concentration in the case of dissolution. For ice melting in salt water, both temperature transport and salt transport contribute to determining melting velocity [11, 12, 13]. This adds a further layer of complexity with opposing contribution of temperature and salt possibly leading to double diffusive convection flow [14], controlling in some conditions the melting velocity [15, 16] and the shape of melting interfaces [17]. Finally, in contrast to the dissolution of salt and sugar, which are denser than the water bath, a buoy is not needed to ensure flotation, as ice floats on water (density $\rho_{ice}=916.7$ $^{-3}<\rho_{water}=999.8 kg m-$3 at the melting temperature $T_{m}=0\,^{\circ}$[18]). The contribution of melting to iceberg drift is currently unknown relative to other contributing factors that include ocean currents, wind, sea slope, surface waves, and Coriolis force [2, 3, 4, 5].
@@
Here, we investigate the kinematics of ice blocks which have asymmetric shapes while floating in a water bath and show that they can not only rotate but translate with significant speeds. We find a typical propulsion velocity of about 3 mm s-1 for triangular ice prisms with an inclined long side of approximately 20 cm and width of approximately 10 cm, floating in a water bath held at a temperature of about $T_{b}=22\,^{\circ}$C, comparable to those observed with dissolving bodies. After presenting the experimental methods in Section II, we demonstrate and quantify the melting-driven propulsion mechanism in Section III. We use shadowgraph imaging [19, 20] to simultaneously track the motion of the block and visualize the buoyancy convection flow. A phenomenological model relating the melting velocity to the terminal speed is developed in Section IV to explain the magnitude of the observed ice block translation velocities as a function of their size, inclination, and bath temperature. In the limit where the heat required to raise the temperature of the ice block to the melting temperature is relatively small, our model finds that although the latent heat plays an important role in the dynamics and determines the time over which the block melts, its actual magnitude does not significantly affect the propulsion speed. In Section IV.4, we compare the measured values over a large range of parameters with the predictions of the model. Then, we demonstrate with experiments that the melting-driven propulsion mechanism extends to baths with ocean water salinity in Section V, and adapt our model to take into account the rising fresh water flow and the falling convection flow due to the cooling of the salt bath. We address the possible relevance of the melting-driven propulsion mechanism to icebergs in oceans in Section VI, before summarizing our main results in Section VII. Although our results cannot be directly applied to icebergs under typical conditions, we argue that this effect may contribute to the motion of icebergs drifting in warmer subpolar regions.
@@
Figure 1: Principle of melting-driven propulsion experiments. (a) A schematic of a melting ice block floating in a water bath. The block is a right-angled triangle prism. The hypotenuse has a length $L$ and is inclined at an angle $\theta$ relatively to the horizontal. The ice block propels along the horizontal coordinate $x$ with a velocity $\bm{U_{b}}$ in steady state. (b) A side view image of a right-angle ice block ($L_{h}=163$ $m$, $W=124$ mm, $H=65$ mm, $\theta=19.5^{\circ}$, $T_{b}=22\,^{\circ}$C). (c) A profile view image of the floating ice block with a parabolic mirror placed behind the tank containing freshwater. The dense cold water leads to convection flow which is visible below the ice block. The mirror is then used for shadowgraph imaging. (d) A shadowgraph picture shows the ice block moving to the right with the convection directed toward the rear of the block (also see Movie S1 [21]).
@@ II Experimental methods
We cast asymmetric ice blocks with a simple model geometry: a right-angled triangle prism. The ice block has a hypotenuse of length $L$ and is inclined at an angle $\theta$ with respect to the horizontal and is shown schematically in Fig. 1(a). The ice blocks are cast in molds with horizontal length $L_{h}$, width $W$, and side $H$ filled with filtered demineralized water and freezing at $15\,^{\circ}$C. Approximately one hundred blocks were cast with various sizes ranging over a few tens of centimeters. The ice blocks contain bubbles which typically results in about 10% of air by volume, similar to icebergs [22]. A few clear rectangular ice blocks which are nearly transparent and defect-free were obtained from the Nice Company (https://www.thenicecompanyparis.com/fr), and cut into triangular prisms with a hot wire. Experiments performed with the clear ice blocks were found to give the same results as the ice blocks with trapped bubbles. The data sets corresponding to various blocks shapes and measurement protocols are listed in Appendix A.
@@
Prior to the commencement of an experiment, an ice block is left to rest at room temperature of about $20\,^{\circ}$C for about ten minutes in order to avoid thermal shock when immersing it into the bath. During this time, the ice temperature approaches the melting temperature of ice. Then, the block is carefully placed with its right angle on the top in a water bath with dimensions that are large compared to the block size. Flotation equilibrium corresponds to the vertical alignment of the gravity center and the center of the immersed part and is typically reached after few oscillations over a few seconds. A side view image of a floating ice block is shown in Fig. 1(b). The actual hypotenuse inclination angle differs from $\theta$ by few degrees due to the buoyancy equilibrium of the asymmetric block (see Appendix B). Yet, we ignore this difference since it appears to have little effect on the measured trends. Ice blocks with $\theta>39.6^{\circ}$ are gravitationally unstable, and we thus focus our study to ice blocks with $\theta$ below this angle.
@@
The ice blocks and their motion in the bath are either observed with a camera from the side using a shadowgraph imaging [19, 20], or from the top. To observe the dynamics with shadowgraphy, we use a glass tank with dimensions $116\times 46$ cm2 filled up to $24$ cm with tap water corresponding to a volume of about $128$ liters. For most experiments the bath temperature $T_{b}$ is the ambient temperature. However, for a few experiments $T_{b}$ has been varied between $10.4$ and $30\,^{\circ}$C using a cooling / heating circulator before performing the measurements. A parabolic mirror with diameter $406$ mm and focal length $1800$ mm is placed on one side of the glass tank at a distance of about 100 mm and oriented parallel to it, as depicted in Fig. 1(c). A small Light Emitting Diode (LED) located at the focus of the parabolic mirror is used for illumination, and a digital camera is located also at the focus of the mirror by the means of a semi-reflective plate. The resulting light beam with nearly parallel rays is refracted by the variations of optical indices due to temperature variations. The projected image captured by the digital camera integrates the density variations along the width of the tank. For a bath of fresh water whose temperature is above maximum density of water at $T_{c}=3.98\,^{\circ}$C, the colder convection plumes sink and and refract ambient light. The plumes appear more clearly in the image shown in Fig. 1(d) generated with shadowgraphy. A small black circle marks the center of the mirror and is visible on some shadowgraph images and in the movies [21]. In these experiments, two 1 mm in diameter nylon wires are positioned just below the surface to guide the motion of the ice block, separated by a distance slightly greater than $W$. This helps maintain the distance between the block and the camera and limits rotation, facilitating observations. In the experiments where the ice blocks are viewed from the top, we use a glass tank with dimensions $90.5\times 44.5$ cm2, filled up to a height of at least $25$ cm with filtered water corresponding to a volume of about $100$ liters. The ice block is free to move in any horizontal direction and data is taken while the block is away from the tank sidewalls. This can be used to study the relative stability of the ice block trajectory.
Figure 2: Observation of a melting pattern on the underside of ice blocks. (a) A view of an ice block floating at the water surface while moving in the direction of the observer after being immersed in the bath for about 8 minutes. A groove pattern generated by melting appears on the bottom inclined face. (b) A schematic cross section of the block illustrating the length-wise grooves. This melting pattern is carved by the backward descending flow. (c) An ice block with bottom surface facing viewer after it is removed from the water bath following approximately 11 minutes of immersion. Typical groove width and depth are about $20$ mm and 55 mm, respectively. This example corresponds to Data Set L. ($L=176$ mm, $\theta=19.5^{\circ}$, $T_{b}=22.1\,^{\circ}$C).
@@ III Experimental demonstration of melting-driven propulsion
Figure 3: Demonstration of the self-propulsion for asymmetric melting ice blocks. (a) Right-angle triangle horizontal ice block position $X_{B}$ as a function of time $t$ plotted along with a linear fit (dashed line) corresponding to a slope $U_{b}=3.02$ mm s-1. Data Set G. (b) The velocity of the block $U(t)$ obtained from $X_{B}$. This experiment corresponds to the block depicted in Movie S2 [21], where the block is kicked manually in the direction opposite to the motion induced by propulsion at $t=8$ s. After a transient, the block recovers the same velocity (dashed line).
@@
To illustrate that the forward motion is quite robust, we manually kicked a block moving at constant velocity in the direction opposite to its motion at approximately 5 s (see Movie S2 [21]). The measured ice block position $X_{B}$, obtained by processing the shadowgraph images, is plotted as a function of time $t$ in Fig. 3(a). The instantaneous velocity $U(t)$ is computed over a moving $1$ s time interval and plotted in Fig. 3(b). The terminal velocity $U_{b}$ is obtained by fitting a line to Fig. 3(a) once steady motion is resumed. We plot a dashed horizontal time in Fig. 3(b) and observe that the block accelerates and reaches the pre-kick velocity $U_{b}\approx 3.02$ mm s-1 in about 30 s. By contrast, symmetric rectangular ice blocks do not show net translation motion (see Movie S3 [21]). We plot $X_{B}(t)$ and $U(t)$ for this symmetric ice block in Fig. 4(a), and Fig. 4(b), respectively. The time scales for motion are significantly larger compared with those observed in the asymmetric examples. The ice block motion shows no clear direction, and capsizing events [27] occur suddenly and modify the convection flow that can change the direction of the ice block drift.
Figure 4: Erratic motion for symmetric melting ice blocks. (a) For a symmetric rectangular ice block position $X_{B}$ shows the slow drift of the block over time (also see Movie S3 [21]). The block dimensions: $100\times 40\times 40$ mm3, $\theta=0^{\circ}$, $T_{b}=20.1\,^{\circ}$C. (b) The corresponding velocity (averaged over $10$ s) shows sudden changes in velocity due to capsizing events at $t=210$ s and $t=506$ s.
@@
We also perform experiments with ice blocks which are unconstrained laterally and we observe their motion in the tank from above. While the sides of the block melt on average at a rate of a few tens of microns per second, we observe that the block inclination is approximately conserved, at least during the first ten minutes or so, even as the ice block shrinks and the edges round over time. However, we find that the movement is not exactly rectilinear, but can show rotation in addition to translation. This rotation may be due to lateral asymmetries in the ice block, and/or caused by the destabilization of the falling convection flow into a vortex as reported with spinning ice disks [8, 28, 29]. The rotational motion becomes increasingly important over time, as the ice block decreases in size and its shape deviates significantly from being prismatic as it completely melts. For this reason, systematic data is taken during the first ten minutes while the overall ice block shape appears similar as the one initially cast, barring the grooves. In these conditions, the dimensions of the blocks considered are determined at the start of the run. The different protocols with various ice blocks shapes are summarized in Appendix A.
@@
The measured terminal velocities $U_{b}$ show more significant experimental variability compared with dissolution propelled bodies [9] that can be attributed to several factors. The density contrast driving the convection flow is stronger by two orders of magnitude in the case of sugar or salt dissolution compared with the density contrast observed in ice blocks when $T_{b}=20\,^{\circ}$C. Consequently, the ice block motion is more sensitive to external perturbations and experimental imperfections. The typical scale of the plumes about $3\delta_{i}\approx 5$ mm as discussed in Section IV.1 is larger compared to the block size leading to greater fluctuations in their trajectories. The shapes of the ice blocks are also less ideal and show around 5% variations. The ice blocks also intermittently release trapped bubbles during melting that further perturb transiently the flow below the ice block. Recent experiments [30] report an enhanced melting rate for vertical ice walls, due to the additional buoyancy associated with the released bubbles. However, in our experiments, we do not find significant differences on average for $U_{b}$ between the standard ice blocks and those made of clear ice without bubbles. Notwithstanding the fluctuations, we observe a robust directed motion with velocity $U_{b}\approx 3$ mm s-1 for an ice block with $L=20$ cm melting in a bath at $2\,{}^{\circ}$C which can be easily noted with the naked eye. The corresponding Reynolds number $Re=\frac{L\,U_{b}}{\nu}$ is of order $600$, with the kinematic viscosity of water $\nu\approx 10^{-6}$ m2 s-1. Thus, the fluid flow is in the inertial regime, and we propose a phenomenological model based on momentum balance inspired by Chaigne, et al. [9].